OpenStudy (anonymous):

Write the equation of the line that passes through (3, 6) and (0, –3) in slope-intercept form. y = 3x – 6 y = 3x – 3 y = 3x + 3 y = 3x + 6

OpenStudy (anonymous):

@jim_thompson5910 can you help

OpenStudy (artsy_brian):

use the slope formula (y2-y1)/x2-x1) -3-6/0-3=-9/-3 which would give the slope equa to 3

jimthompson5910 (jim_thompson5910):

another way is to plug each point into each equation

jimthompson5910 (jim_thompson5910):

if both points satisfy the equation, then you found your answer

OpenStudy (anonymous):

wait @jim_thompson5910 can you show me what i have to plug in for the equation

jimthompson5910 (jim_thompson5910):

if you look at choice A) y = 3x – 6

jimthompson5910 (jim_thompson5910):

and plug in (3,6) you get this

OpenStudy (anonymous):

alright i will do that now

jimthompson5910 (jim_thompson5910):

y = 3x – 6 6 = 3(3) – 6 ... plug in x = 3 and y = 6 6 = 9 - 6 6 = 3 since that equation is false, this means that (3,6) is NOT on the line y = 3x – 6

OpenStudy (anonymous):

what about 0,-6 what happened to that?

jimthompson5910 (jim_thompson5910):

0,-6 ?

OpenStudy (anonymous):

-3*

jimthompson5910 (jim_thompson5910):

well if the first point does NOT work for y = 3x – 6, then what does it matter if the second point works or not

jimthompson5910 (jim_thompson5910):

BOTH points need to work

OpenStudy (anonymous):

for choice B is works out

jimthompson5910 (jim_thompson5910):

so if one doesn't work, then you can eliminate that answer choice

jimthompson5910 (jim_thompson5910):

yep it's choice B